OpenStudy (anonymous):

Volume of a circle equation? Calculate the volume of liquid in the round flask sketched below. Give your answer in liters, and round to the nearest . the diameter of the flask is 22.0 cm

Directrix (directrix):

If this is a volume of a sphere problem, then the sphere formula is V = 4/3 * π r³. r = 22/2 cm.